OpenStudy (i_love_my_nieces):

A bank offers two interest account plans. Plan A gives you 6% interest compound annually. Plan B gives you 13% annual simple interest. You have to invest $2,000 for the next 4 years. Which account earns you the more interest (in dollars) after 4 years? How much will you have earned?

OpenStudy (i_love_my_nieces):

@SolomonZelman

OpenStudy (i_love_my_nieces):

@SolomonZelman can you please help me?

OpenStudy (solomonzelman):

\(\huge\color{blue}{ A=P(1+\frac{r}{n} )^{nt}}\) P = principal amount (the initial amount you borrow or deposit) r = annual rate of interest (as a decimal) t = number of years the amount is deposited or borrowed for. A = amount of money accumulated after n years, including interest. n = number of times the interest is compounded per year \(\huge\color{blue}{ A=2000(1+\frac{.06}{1} )^{(1 \times 4)}=?}\) (you tell me) this is compound annually at 6%

OpenStudy (i_love_my_nieces):

Idk.

OpenStudy (solomonzelman):

you can't calculate this ?

OpenStudy (i_love_my_nieces):

I don't have a calculator to do it :'(

OpenStudy (solomonzelman):

the other one 13% simple interest. \(\huge\color{blue}{ 2000 \times 0.13 }\) <-- the interset you get each year. So that times 4 + the original 2000.

OpenStudy (solomonzelman):

can you at least do the simple annual 13% interest for me?

OpenStudy (i_love_my_nieces):

1040?

OpenStudy (solomonzelman):

that's what you get each year. There are 4 years, so the interest altogether would be 260*4 e.i. 1040 + the principal amount (the deposit) 2000+1040 ... = 3040

OpenStudy (solomonzelman):

now, .... annually compounded at 6% interest.